OpenStudy (anonymous):

A six-sided number cube with faces numbered 1, 2, 3, 4, 5, and 6 is rolled. What is the probability of rolling a 2 or a number greater than 5? A. 1/36 B. 1/3 C. 1/2 D. 2/3

OpenStudy (anonymous):

Would this be 2/3?

OpenStudy (anonymous):

Or 1/3?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Circle the numbers that are either... a) equal to 2 OR b) greater than 5

OpenStudy (anonymous):

So 2 and 6

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

there are 2 numbers (2 and 6) in the event space there are 6 numbers (1,2,3,4,5,6) in the sample space so... P(A) = 2/6 = 1/3 where A = event that you roll a 2 or a number greater than 5

OpenStudy (anonymous):

Okay tysm, I understand now

jimthompson5910 (jim_thompson5910):

I'm glad it's making sense